OpenStudy (anonymous):

7b^2 + 9b + 2b^2 + 8 How do I solve this??

OpenStudy (anonymous):

add 7b^2 and 2b^2

OpenStudy (anonymous):

9b^4?

OpenStudy (anonymous):

No. You have 7 * b^2 and you also have 2 * b^2 How many "b^2"s is that?

OpenStudy (skullpatrol):

$$7b^2 + 9b + 2b^2 + 8$$

OpenStudy (anonymous):

thats 2 b^2's

OpenStudy (skullpatrol):

$$7b*b + 9b + 2b*b + 8$$

OpenStudy (anonymous):

NO!!! "b^2" is a quantity You have 7 of this quantity and you also have 2 of this quantity. How many of these quantities do you have if you add them together?

OpenStudy (skullpatrol):

$$7b^2 + 9b + 2b^2 + 8$$ $$= 7b^2 + 2b^2 +9b+ 8$$

OpenStudy (skullpatrol):

$$(7+2)b^2 +9b+8$$

OpenStudy (anonymous):

Ah I see now, what if it were 2^2 + 3^2? Would it be 5^2?

OpenStudy (skullpatrol):

$$2^2=2*2$$ $$3^2=3*3$$

OpenStudy (anonymous):

No. Exponents don't work that way. Look. "b^2" is a quantity like "X" How many Xs do you have if you add 7 Xs and 2 Xs together?

OpenStudy (skullpatrol):

$$2^2 + 3^2 =?= 5^2$$ $$2*2+3*3 =?= 5*5$$ $$4 + 9 =?= 25$$ $$13 \neq 15$$

OpenStudy (anonymous):

How much is 7x + 2x?

OpenStudy (anonymous):

9x

OpenStudy (anonymous):

Exactly! So then how much is 7b^2 + 2b^2?

OpenStudy (anonymous):

9b^2

OpenStudy (anonymous):

Correct!!! :-) Now rewrite your formula using these combined terms.